# Decoding Intelligence: A Framework for Certifying Knowledge Comprehension in LLMs

### v2

@@ 4.3 Certificates
QuaCer-C generates certificates providing high-confidence, tight lower and upper bounds on the probability of a correct LLM response to a random prompt sampled using the prompt constructor from the given distribution of prompts in our specifications. We report the average value of the lower and upper bounds, over our set of specifications that QuaCer-C certifies for each LLM, in Table 1. We also report the average empirical probability (the ratio of correct responses to the total number of prompts, $n$, for each certificate), averaged over the test set. Next, we summarize the key observations from the certification results in Table 1, some of which follow trends from prior works.
Scaling of knowledge comprehension with model size. We observe that the larger models such as Gemini and GPT have significantly higher bounds than those for the smaller models such as Phi-3, Mistral, Llama (Figure 3).
Figure 3: Variation in certification bounds with models (Vanilla, fp16)
@@
The lower bounds of the larger models are higher than the upper bounds of the smaller models, suggesting a paradigm shift in knowledge comprehension capabilities (especially for Gemini-1.5-Pro and GPT-4o). However, as the larger models are also closed-source, we are unaware whether the enhanced knowledge comprehension capabilities could be due to specialized training or finetuning techniques applied on the models. We see that the smaller models have similar certification bounds. Interestingly, Phi3-3B, which is the smallest model we consider, is performing comparatively to its 14B counterparts and also to larger Mistral and Llama models.
@@
Effects of model quantization. We see that with increase in quantization, the model performance deteriorates on knowledge comprehension, which is contrary to prior findings such as Jin et al. (2024) that suggest that 4-bit quantization can retain the model’s knowledge and reasoning capabilities. The drop in the performance is not large, however.
Effects of different kinds of specifications. Our results for the different kinds of specifications — Vanilla, Shuffle, and Shuffle Distractor, indicate that the Vanilla specifications are generally easier for models, leading to higher certification bounds than those for the other specifications. Shuffle Distractor specifications are challenging specifications for all models resulting in consistently lower certification bounds. However, the differences in the bounds’ values are not high across the settings, which could be because of the challenge of large and unstructured contexts from which the model need to identify relevant information in all cases.
Quality of bounding intervals. Table 1 presents the average bounds from the confidence intervals generated for each certificate. A desirable property for the intervals, alongside their high confidence, is that they should be tight, i.e., their range should be small. Tighter intervals indicate precise analysis with small errors. We observe that the average range of the confidence intervals obtained in our experiments is less than $0.12$.
@@ 4.4 Case studies
Figure 4: Qualitative analysis of the samples used for certifying knowledge comprehension for the Vanilla specifications on the Wikidata5m subgraph pivoted at the node for ‘Batman Begins’ movie. The context provided in the prompts is not shown for brevity. Wrong model responses are colored red and the correct ones are colored green. The samples are consistent with the certification results, wherein Phi-3 (3B) has lower certification bounds than those for GPT-4o, which has lower bounds than those for Gemini-Pro.
Figure 5: Categorizing predominant types of LLM responses. The LLMs can either do good reasoning or show failure instances caused by distractors or missed reasoning steps.
In this section, we conduct a qualitative analysis of the certification results in Table 1. Firstly, we show the responses of $3$ models in Figure 4 — Phi-3 (3B), GPT-4o, and Gemini — that were obtained when certifying them for the Vanilla specification defined over a subgraph pivoted at the node for ‘Batman Begins’ movie. The samples are representative of the models’ certification results.
Next, we categorize and identify the predominant kinds of model responses and present their examples. In particular, we frequently see the following two failure modes — distracted and missed relation. In the former case, the model gets deviated from the query by following the distractor context in its prompt, resulting in an incorrect answer. In the latter, the model skips some reasoning steps to arrive at the final, correct answer. In other cases of good reasoning, the model accurately follows the steps of the query and arrives at the correct answer. Figure 5 presents examples of the aforementioned kinds of model responses for the GPT-4o model.
@@ 5 Related works
In-context learning. As the context windows of LLMs increase (Gemini Team, 2024; Chen et al., 2023; Dubey et al., 2024), more information can be provided to the models to improve their performance. Moreover, few-shot demonstrations (Brown et al., 2020) and examples from related tasks (Qu et al., 2024) can be provided in the context to inform the models about task-specific details, which can result in better performance. This emergent behavior (Wei et al., 2022b; Lu et al., 2024) wherein LLMs can become proficient at a task with few demonstrations provided in their prompts is in-context learning. We use in-context learning to enhance LLMs’ knowledge to answer the prompts and provide few-shot demonstrations for the reasoning steps expected from the models.
@@
Benchmarking LLM intelligence. Several benchmarks have been proposed to study the reasoning (Zhou et al., 2022; Huang & Chang, 2023; Plaat et al., 2024), arithmetic (Yuan et al., 2023; Song et al., 2024; Yang et al., 2023a), planning (Pallagani et al., 2023; Valmeekam et al., 2023; Huang et al., 2022), and question-answering (Yang et al., 2018; Ho et al., 2020; Welbl et al., 2018) capabilities of LLMs, which are integral components of human intelligence. These benchmarks provide empirical insights and trends into the performance of LMs. However, these insights are generally for static datasets and are not guaranteed to generalize. On the other hand, certification methods provide guarantees on, for example, the scope (defined by specifications) and confidence of its claims, as we illustrate in this work.
6 Conclusion and future work
In this work, we present a novel framework to formally certify LLMs for knowledge comprehension. We develop novel specifications that quantify the probability of correct responses over any random knowledge comprehension prompts from a distribution derived from a knowledge graph. Certificates for knowledge comprehension consist of high-confidence bounds on the probability of correct knowledge comprehension, thus providing a method to compare different LLMs with formal guarantees. Our experiments study the variations in knowledge comprehension along the axes of model size, quantization, and task difficulty and present novel trends.
As our work is the first to certify knowledge comprehension in LLMs to the best of our knowledge, there are several possible directions for future work. Our framework can be integrated with knowledge graph construction methods (Ye et al., 2022), to specify and certify LLMs for comprehension and reasoning over less structured/proprietary inputs. The certification method can also be made more sample-efficient. Moreover, our framework can be adapted to domains with other kinds of graphs, such as social network graphs or medical graphs, to certify LLMs for comprehension in domains of social importance.
Appendix A Ablations
A.1 Few Shot Prompts
Table 2: Certification results for different LLMs with different number of few-shot examples
Model Avg. lower bound Avg. upper bound Avg. accuracy
Gemini-1.5-Flash Vanilla (Default) $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
Gemini-1.5-Flash 3Shot Vanilla $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
Gemini-1.5-Flash 4Shot Vanilla $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
Gemini-1.5-Flash 5Shot Vanilla $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
A.2 Distractor Distributions
  1. 1.
  2. 2.
  3. 3.
Model Avg. lower bound Avg. upper bound Avg. accuracy
Gemini-1.5-Flash Setting 2 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
Gemini-1.5-Flash Setting 3 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
A.3 Model Performances with varying Path Length
Varying Setting: In figure  6 we show plots for various specifications for the GPT4o model.
Varying Quantization: In figure  7 we show plots when the quantization is varied with the Llama3-8B model on the shuffle specification and their effects on performance.
Varying Models: In fig  8 we show plots for the shuffle specification and performance across the models(the open-source models use fp16 precision).
In Figure  6, we observe that the performance across settings converges as k increases and the distractor setting is most impactful on the performance for $k=2$.
In Figure  7, we infer that as k increases the performance of the models’ on the task converges across the different quantizations. We hypothesize this is due to the increasing complexity of the reasoning task.
Appendix B Knowledge Graph and Query Generation
@@
B.1 Knowledge Graph Structure
Our experiments are based on the Wikidata5m knowledge graph (KG). The KG has the following key characteristics:
  • Nodes: Each node represents an entity and is associated with a text paragraph from Wikidata5m.
  • Edges: Edges represent relationships between entities.
  • Text Paragraphs: The text associated with each node may contain information relevant to its connected edges.
  • Node and Edge Aliases: Each node and each edge has a set of aliases associated with them which are just different names for them.
This structure allows us to generate queries that require reasoning across multiple hops in the graph.
B.1.1 Preprocessing the wikidata5m knowledge graph
@@
B.2 Query Generation
We utilize the Wikidata5m knowledge graph for multi-hop query generation. The query generation process involves the following steps:
B.2.1 Random Path Generation
We begin by selecting a pivot node $v_{0}$ in the knowledge graph ${\mathcal{G}}$. From this pivot, we construct a local subgraph ${\mathcal{G}}(v_{0})$ consisting of all paths $\Pi_{v_{0}}$ originating from $v$. This local subgraph serves as the domain for our path generation process. As arbitrary long paths can lose their semantic meaningfulness, we use a constraint $\rho$ to restrict the length of paths from the pivot node in the subgraph to be maximum $\rho$.
Within ${\mathcal{G}}(v_{0})$, we generate a path $\Pi$ using a randomized depth-first search algorithm. The length of this path, denoted as $k_{choice}$, is sampled randomly from the set ${1,2,...,\rho}$ according to a discrete uniform distribution.
This randomized depth-first search traverses the neighbors of each node in ${\mathcal{G}}(v_{0})$ in a random order, which directly corresponds to the sampling process described in line 10 of Algorithm 1. Specifically, at each step, we sample the next node in the path from a discrete uniform distribution over the current node’s neighbors within the local subgraph, expressed as $\sim(\mathcal{D}([v^{\prime}\mid(v_{i},v^{\prime})\in{\mathcal{E}}\land v^{\prime}\in{\mathcal{G}}(v_{0})]))$, where $v_{i}$ is the current node in the path.
@@ B.2.1 Random Path Generation
To ensure well-defined queries with unique answers, we introduce an additional constraint on path generation. This constraint requires that each generated path be unique in terms of its sequence of relationships. Specifically, traversing the path from the initial node using the specified relations must lead to a single, unambiguous answer node. This approach prevents queries with multiple valid answers, which would complicate the evaluation of the language model’s performance. It’s important to note that this uniqueness constraint applies only to the specific path being generated. Nodes within the path may still have multiple edges with the same relation type to other entities not on the path. This allowance maintains the natural complexity of the knowledge graph structure, where entities can have multiple relationships of the same type with different entities.
The pseudocode for the path generation algorithm is specified in  2
@@
B.2.2 Query Formulation
Once a valid path $\Pi$ is generated, we convert it into a query string. This process aligns with line 11 in Algorithm 1. The query is constructed by sampling aliases for each node and relation in the path. For example, a path $\Pi=[A,B,C]$ might be converted to a query ”sampled_alias(A) $\rightarrow$ sampled_alias((A, B)) $\rightarrow$ sampled_alias(B) $\rightarrow$ sampled_alias((B, C)) $\rightarrow$ ?”. Here the tuple of two nodes represents their edge. The aliases are sampled randomly from a discrete uniform distribution over the available aliases for a node or an edge.
B.2.3 Example Query Generation
To illustrate our query generation process, consider the scenario of a path in our subgraph as shown in  9.
@@ B.2.3 Example Query Generation
Our algorithm would construct the following query from the path presented in  9:
@@
B.3 Prompt Construction

### v3

@@ 3.1 Certificates
LLMCert-C generates certificates with high confidence ($1-\delta=95\%$), tight lower and upper bounds on the probability of a correct LLM response to a random prompt sampled from the prompt distribution in our specifications. LLMCert-C uses $n=250$ number of samples. We report the average value of the lower and upper bounds over all specifications over a KG that LLMCert-C certifies for each LLM, in Table 1. Each certificate takes 10-20 minutes for generation, where the main bottleneck is each LLM inference. We plan to optimize this in future work. We also report the average baseline result for each static dataset from the specifications. Appendix E presents detailed ablation studies for the impact of the various certification hyperparameters. Next, we summarize the key observations from Table 1.
Figure 2: Samples from PrimeKG’s Distractor specifications. Prompt context not shown for brevity. Wrong/correct responses are colored red/green respectively. Samples are consistent with certificates, where Phi-3 (14B) has lower bounds than GPT-4o and Gemini-Pro.
Figure 3: Negative effects of distractors on LLM performance for answering queries from same DAGs (correct and wrong answers in green and red respectively).
@@
Variation in knowledge comprehension with models. We observe that larger models, such as Gemini and GPT, have significantly higher bounds than those for the smaller models, such as Phi-3, Llama, for specifications over both KGs. The lower bounds of the larger models are higher than the upper bounds of the smaller models, suggesting a paradigm shift in knowledge comprehension capabilities (especially for Gemini-1.5-Pro and GPT-4o). However, as the larger models are also closed-source, we are unaware whether the enhanced knowledge comprehension capabilities could be due to specialized training or fine-tuning techniques applied on the models. From the certification bounds for PrimeKG, we can establish performance hierarchies with high confidence across the evaluated models. Specifically, when we observe higher lower bound of a model than the upper bound of another model (e.g., Gemini-v1.5-Pro vs Llama 11B for all settings), we conclude the superior performance of the former in the worst case with high confidence.
Effects of different kinds of specifications. Our results for the different kinds of specifications indicate that Vanilla specifications are easier and lead to higher certification bounds. Distractor specifications are challenging for all models, leading to upper bounds that are consistently lower than the lower bounds for the Vanilla setting. This suggests susceptibility of SOTA LLMs to natural noise in prompts. Even larger models such as Gemini-1.5-Pro, which show close to perfect performance in the baseline and Vanilla case (for PrimeKG only), can be incorrect with distractors, showing the nontrivial effects of noise.
Comparison with benchmarking baseline. Baseline scores of all models approach or surpass average certification upper bounds, suggesting inflation of performance estimates in benchmarking. For example, contrary to certification bounds for PrimeKG’s distractor specifications, Phi-3 (14B) outperformed Llama (11B). Such findings emphasize the need for more reliable evaluation methods grounded with statistical guarantees. Our certificates also improve upon the baseline by preventing incorrect comparisons between models with similar performance (e.g., Phi-3 14B and Llama 11B and respectively between their 4-bit models for Vanilla specifications) by showing overlaps in their certification bounds.
@@
Effects of model quantization. We see that higher quantization deteriorates model performance on knowledge comprehension, contrary to prior works like Jin et al. (2024) that suggest that 4-bit quantization can retain the model’s knowledge and reasoning capabilities.
@@ 3.2 Case studies
Next, we qualitatively analyze certificates for PrimeKG specifications. First, we show the responses of $3$ models in Figure 2 — Phi-3 (3B), GPT-4o, and Gemini-Pro, obtained when certifying them for the Distractor specification. The samples reflect the certification bounds, showing the performance hierarchy across the models. The certification bounds can be used to select the best models for critical applications. Next, we identify the vulnerabilities of SOTA LLMs to distractor texts in Figure 3. We see consistent occurrences of incorrect model responses in the presence of distractors and not otherwise. This demonstrates the vulnerabilities of SOTA LLMs and the risks posed by their deployment in critical applications. We discuss similar case studies for Wikidata5m specifications in Appendix F.
@@ 4 Related works
In-context learning. As LLM context windows increase (Gemini Team, 2024; Chen et al., 2023; Dubey et al., 2024), more information can be provided in the prompts like few-shot demonstrations (Brown et al., 2020) and examples from related tasks (Qu et al., 2024). In-context learning is the emergent behavior (Wei et al., 2022; Lu et al., 2024) in which LLMs become proficient at a task with demonstrations in prompts. We use in-context learning and few-shot examples to enhance LLMs’ knowledge comprehension.
@@
Benchmarking LLM intelligence. Several benchmarks have been proposed to study reasoning (Huang & Chang, 2023; Plaat et al., 2024; He et al., 2024; Zha et al., 2021), arithmetic (Yuan et al., 2023; Song et al., 2024), planning (Pallagani et al., 2023; Valmeekam et al., 2023), and question-answering (Yang et al., 2018; Welbl et al., 2018; Perevalov et al., 2022) in LLMs. These benchmarks provide empirical insights and trends of LM performance. However, they use static datasets and not guaranteed to generalize. Certification provides guarantees on the scope (defined by specifications) and confidence of the claims, as shown in this work.
Appendix A A note for practitioners
Our work takes the first step towards providing formal guarantees of LLMs’ knowledge comprehension capabilities, with particular relevance for high-stakes domains like healthcare. While our framework enables more reliable LLM deployment through rigorous guarantees, these certificates should be viewed as probabilistic bounds rather than absolute guarantees of model reliability. We anticipate this work will help make LLMs more trustworthy for safety-critical applications, although continued human supervision remains essential. Future work can extend our framework to certify additional properties like safety and robustness, and integrate it with knowledge graph construction methods to handle less structured inputs. With a one-time effort to construct domain-specific knowledge graphs, practitioners can obtain reliable insights into the performance of LLMs for question-answering and knowledge comprehension tasks. We release our implementation to support further research on formal verification of language models.
Appendix B Background
B.1 Knowledge graph
A knowledge graph ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ is a collection of nodes ${\mathcal{N}}$ representing entities, interconnected by directed edges ${\mathcal{E}}$ representing their relations (Peng et al., 2023; Ji et al., 2022). They are commonly used in search engines to improve the accuracy of responses to user queries. Hence, big companies develop their own closed-source knowledge graphs. PrimeKG (Chandak et al., 2023) is an open-source knowledge graph of precision medicine consisting of nodes corresponding to diseases, drugs, effects, etc. PrimeKG describes $\sim$17k diseases with $\sim$4M relations. Wikidata5m (Wang et al., 2021b), another popular open-source knowledge graph consisting of 5M nodes, is a structured representation of Wikipedia. Each Wikidata node corresponds to a Wikipedia page, containing its abstract and a set of aliases that can synonymously refer to the node. Two nodes $(v_{1},v_{2})$, $v_{1},v_{2}\in{\mathcal{N}}$ are connected by a labeled edge if there is a link in the supporting document for $v_{1}$ to that for $v_{2}$. Edge $(v_{1},v_{2})$ is labeled by a set of aliases for the relation between $v_{1}$ and $v_{2}$. App. Figure 4 shows a subgraph of Wikidata5m.
Figure 4: A subgraph of Wikidata5m
B.2 Large Language Models
Large Language Models (LLMs) are autoregressive causal language models that operate on a vocabulary $\mathcal{V}$, a set of tokens. LLMs takes a sequence of tokens $x_{1},...,x_{n}$ where $x_{i}\in\mathcal{V},n>0$ and outputs a probability distribution over $\mathcal{V}$ to sample the next token $x_{n+1}$. These models are pretrained on vast corpora (Liu et al., 2024) and have shown remarkable capabilities  (Touvron et al., 2023; Gemini Team, 2024; OpenAI, 2024). Numerous benchmarks (Yang et al., 2018; Rein et al., 2023; Hendrycks et al., 2021) have been developed to evaluate the performance of LLMs on tasks related to multi-step reasoning, knowledge comprehension and question answering. However, there remains a gap in our theoretical understanding of LLMs’ capabilities.
Appendix C Validity of Certification bounds
In this section, we provide empirical evidence for the validity of the certification bounds with the user-defined confidence (95%). We select $10$ equally-spaced values of the true probability of correct LLM response, between 0 and 1. For each true probability, we generate $1000$ Clopper-Pearson confidence intervals, which are used to form our certification bounds. We calculate the proportion of the confidence intervals that contain the true probability and compare with the confidence level. We show in Figure 5 that the proportion of instances where the confidence intervals contain the true probability is almost always more than $95\%$.
Figure 5: Variation in the proportion of Clopper-Pearson confidence intervals containing the true probability of correct response
Appendix D Knowledge Graph and Query Generation
@@
D.1 Knowledge Graph Structure
Our experiments utilize two knowledge graphs: Wikidata5m and PrimeKG, each with distinct characteristics.
D.2 Wikidata5m
Wikidata5m is characterized by rich textual and semantic features:
  • Nodes: Each node represents an entity and is accompanied by a descriptive text paragraph that provides contextual information about the entity.
  • Edges: Relationships between entities are represented as edges.
  • Text Paragraphs: Each node contains an associated text paragraph that that often provides at minimum information relevant to its relationships with other entities.
  • Aliases: Each node and each edge has a set of aliases associated with them which are just different names for them.
D.2.1 Preprocessing the wikidata5m knowledge graph
@@
D.3 PrimeKG
PrimeKG presents a more streamlined structure:
  • Nodes: The graph consists of nodes representing entities, but unlike Wikidata5m, nodes do not contain accompanying descriptive text.
  • Edges: Similar to Wikidata5m, edges denote relationships between entities.
  • Aliases: Each node and edge is associated with exactly one alias, providing a single alternative name rather than multiple variants.
D.3.1 Preprocessing the PrimeKG knowledge graph
The PrimeKG dataset requires several key transformations to align with our experimental framework.
  1. 1.
    Converting to a Directed Graph: We convert PrimeKG’s undirected structure to a directed graph by creating bidirectional edges. Each original edge generates two directed edges with labels incorporating the target entity’s type (e.g., an undirected indication edge between a drug and disease yields ”indication (drug)” and ”indication (disease)” directed edges).
  2. 2.
    Enriching Entity Representations:We augment entity representations by prepending type information to each entity’s alias, maintaining a single alias per node to preserve medical precision.
  3. 3.
    Generating Supporting Text: To address PrimeKG’s lack of contextual information, we generate descriptive text for each edge using templates tailored to the entity types and relationships (e.g., ”entity_1 is relation name for entity_2”). Implementation details are available in the accompanying code.
  4. 4.
    Standardizing Character Encoding: All text is converted to ASCII format, matching the Wikidata5m preprocessing pipeline.
D.4 Query Generation
Query generation follows a two-phase process: (1) DAG generation according to the knowledge graph structure and specifications, and (2) transformation of the DAG into a natural language query. The specifics differ between Wikidata5m and PrimeKG due to their distinct structures and certification requirements.
D.4.1 Wikidata5m
For Wikidata5m, we generate linear DAGs (paths) from specified pivot nodes. This approach enables focused testing of an LLM’s knowledge comprehension capabilities through a chain of relationships. In this case therefore ${\mathcal{N}}_{source}$ has only one node the pivot node, $v_{0}$. Then we simulate the algorithm 2 and sample a path for our query generation.
From the pivot node $v_{0}$, we construct a local subgraph ${\mathcal{G}}(v_{0})$ containing all paths ${\mathcal{N}}_{v_{0}}$ originating from $v_{0}$. This local subgraph construction serves two purposes: efficient path sampling and enforcement of the $\rho$ constraint from Algorithm 1. We set $\rho=5$ as longer paths often lose semantic coherence in knowledge comprehension queries.
Within ${\mathcal{G}}(v_{0})$, we generate paths through a randomized sampling process:
1. Sample path length $k_{choice}$ uniformly from $\{1,\ldots,\rho\}$, following line 1 in Algorithm 2
2. Generate path through iterative node sampling: - Start with ${\mathcal{N}}=[v_{0}]$ At each step $i$, sample next node according to distribution $\mathcal{D}$ (defined in line 1):
$v_{i+1}\sim\mathcal{D}([v^{\prime}\mid(v_{i},v^{\prime})\in{\mathcal{E}}\land v^{\prime}\in{\mathcal{G}}(v_{0})])$
where $v_{i}$ is the current terminal node of ${\mathcal{N}}$
@@ 1. Sample path length $k_{choice}$ uniformly from $\{1,\ldots,\rho\}$, following line 1 in Algorithm 2
To ensure well-defined queries per Definition 2.1, we require path uniqueness: following the relation sequence of ${\mathcal{N}}$ from any node must lead to at most one target node. This constraint ensures $|{\mathcal{N}}^{\prime}_{sink}|=1$ as specified in Definition 2.1, while preserving the natural graph structure where nodes may have multiple edges with the same relation to non-path nodes. Additionally, this approach prevents queries with multiple valid answers, which would complicate the evaluation of the language model’s performance. It’s important to note that this uniqueness constraint applies only to the specific path being generated. Nodes within the path may still have multiple edges with the same relation type to other entities not on the path. This allowance maintains the natural complexity of the knowledge graph structure, where entities can have multiple relationships of the same type with different entities.
The complete path generation process is detailed in Algorithm 4, which implements the path sampling component of sampleDAG used in lines 2-3 of Algorithm 2.
The pseudocode for the path generation algorithm is specified in  4
@@
Query Formulation:
Once a valid path ${\mathcal{N}}$ is generated, we convert it into a query string. This process aligns with line 3 in Algorithm 1. The query is constructed by sampling aliases for each node and relation in the path. For example, a path ${\mathcal{N}}=[A,B,C]$ might be converted to a query ”sampled_alias(A) $\rightarrow$ sampled_alias((A, B)) $\rightarrow$ sampled_alias(B) $\rightarrow$ sampled_alias((B, C)) $\rightarrow$ ?”. Here the tuple of two nodes represents their edge. The aliases are sampled randomly from a discrete uniform distribution over the available aliases for a node or an edge.
Example Query Generation
To illustrate our query generation process, consider the scenario of a path in our subgraph as shown in  6.
@@
Our algorithm would construct the following query from the path presented in  6:
@@
D.4.2 PrimeKG
For PrimeKG, each certificate is defined by a reference DAG ${\mathcal{G}}_{ref}=({\mathcal{N}}_{ref},{\mathcal{E}}_{ref})$ that encodes a specific clinical reasoning pattern. Following Algorithm 3, we generate isomorphic DAGs ${\mathcal{G}}^{\prime}$ that preserve the structure and relation types of ${\mathcal{G}}_{ref}$ while varying the specific entities.
Formally, for a given ${\mathcal{G}}_{ref}$, we sample DAGs ${\mathcal{G}}^{\prime}$ such that there exists a bijection $\phi:{\mathcal{N}}_{ref}\rightarrow{\mathcal{N}}^{\prime}$ preserving: 1. Edge structure: $(v_{1},v_{2})\in{\mathcal{E}}_{ref}\iff(\phi(v_{1}),\phi(v_{2}))\in{\mathcal{E}}^{\prime}$ 2. Relation types: $(v_{1},v_{2})_{\mathcal{A}}=(\phi(v_{1}),\phi(v_{2}))_{\mathcal{A}}$
We define nine core reference patterns capturing common clinical reasoning chains:
  • Which drug is a contraindication for #disease A# and indication for #disease B#, when patient has both diseases?
  • Which drug can be used as indication for ¡disease A¿ and is atleast an off-label use drug for #disease B#?
  • Which drug can be used as an indication for #disease A# and is an indication for #disease B#?
  • Which drug is an indication for #disease A# and interacts synergistically with the treatment of #disease B#?
  • Which drug targets #genes/proteins(RIDR33)# associated with #disease(RIDR38)#?
  • Which disease is contraindicated by drug indication #drug A# for #disease A#, which both show #effect/phenotype A#?
  • Which disease is treated with #drug A# but contraindicated with #drug B#?
  • Which drug can be used to treat a disease caused by exposure of #exposure A#?
  • Which disease can be a diagnosis for the following phenotypes - #list of phenotypes#?
For each reference DAG ${\mathcal{G}}_{ref}$, we implement specialized sampling functions that generate isomorphic DAGs while respecting PrimeKG’s medical domain constraints. These functions implement the generateIsomorphisms procedure from Algorithm 3 as follows:
1. Type constraints: Each node $v\in{\mathcal{N}}_{ref}$ specifies required entity types (e.g., DRUG, DISEASE) 2. Relation constraints: Each edge $e\in{\mathcal{E}}_{ref}$ requires specific medical relationships (e.g., indication, contraindication) 3. Clinical validity: Additional constraints ensure sampled DAGs represent valid clinical scenarios
For implementation details of these specialized sampling functions, see the discovery_functions in our codebase.
Query Formulation:
For a sampled DAG ${\mathcal{G}}^{\prime}$, we generate natural language queries using predefined templates $\tau$ (line 4 in Algorithm 1). Each certificate has a minimum of 5 semantically equivalent templates that vary in:
1. Syntactic structure: ”Which drug treats X and Y?” vs. ”Find a medication indicated for both X and Y” 2. Clinical terminology: ”contraindication” vs. ”should not be used with” 3. Question framing: Direct vs. scenario-based queries
D.5 Prompt Construction